# Bifurcations of Highly Inclined Near Halo Orbits using Moser Regularization

### v1

@@ Abstract.
1. Introduction
@@
A detailed numerical study of near-collision periodic orbits requires a regularization framework compatible with high-precision continuation. The Kustaanheimo-Stiefel (KS) regularization [Kustaanheimo_stiefel_regularization_1965], as described for the numerical setting in [howell_breakwell_almost_rectilinear_halo_1984], has traditionally been used for this purpose. In this study, we adopt an alternative approach based on Moser’s regularization. This framework offers several advantages: it is applicable in arbitrary dimensions (e.g. planar or spatial CR3BP), preserves periodicity of orbits between the regularized and unregularized systems, and provides a direct correspondence of Floquet multipliers (Lemma 3.4), simplifying stability analysis. We provide a first, complete implementation of Moser regularization for numerical continuation and stability analysis of periodic orbits. Our implementation treats the dynamics as a Hamiltonian system with constraints, whose framework we review in Appendix B. In particular, the regularized phase space appears as a $6$-dimensional constrained submanifold embedded in an $8$-dimensional Euclidean space.
@@
Table 1. Summary of the orbit families studied, together with their symmetries and Conley-Zehnder indices. Symmetric orbits are symmetric with respect to $y=0$, while those marked with $r_{x}$ are symmetric with respect to $x=0$.
@@
$L_{1}$ halo (collisional) small $\mu>0$ symmetric 3, 2 4.2 and 4.3
@@
Saturn-Enceladus symmetric 10, 11, 12
tri-fly (“up”, “down”) Hill’s problem symmetric ($r_{x}$) 10, 11
Saturn-Enceladus non-symmetric 10, 11
W4/W5 Hill’s problem symmetric ($r_{x}$) 3 4.5
@@
For detailed analysis, we compute the Conley-Zehnder indices of the periodic orbit families in this study; the Conley-Zehnder index is a symplectic invariant that remains constant along non-degenerate continuous families and whose change implies bifurcations. This index provides both a diagnostic tool for numerical detection of bifurcations and a topological classification of families, and has been recently used for numerical bifurcation studies of CR3BP and related systems [moreno_aydin_van_Koert_Frauenfelder_Koh_Bifurcation_Graphs_2024, aydin_Batkhin_hill_2025, aydin_dro_2025]. We give details and computation methods in Appendix D.
The paper is organized as follows. The main text focuses on the underlying celestial mechanics, while the contents of a purely mathematical nature is collected in the appendix. A brief outline is given below. Section 2 introduces the rescaled Hamiltonian formulation for CR3BP and its limiting relation to Hill’s problem. Section 3 presents the numerical framework for Moser regularization and introduces the vertical collision orbit. Section 4 presents the numerical results. In Appendix B, we provide details on Hamiltonians with constraints and the restricted phase space. We follow that in Appendix C with additional details of differential correction in Moser regularization. Appendix D discusses the Conley-Zehnder index and computations of symplectic invariants such as local Floer homology.
2. Dynamical Models and Symplectic Scaling Framework
2.1. The Circular Restricted Three-Body Problem (CR3BP)
@@
Table 2. Table of mass ratio parameters for different three-body systems.
2.2. Hill’s Problem and Symmetries
@@
Figure 1. Configuration of the Lagrange points and reflection symmetries in the CR3BP (left) and Hill’s problem (right). In both models, the reflection symmetry $r_{y}$ across the plane $y=0$ is present, while the Hill’s problem exhibits additional symmetry $r_{x}$ across the plane $x=0$. Only $L_{1}$ and $L_{2}$ Lagrange points remain in Hill’s problem.
2.3. Rescaled CR3BP and the Hill’s Limit
@@
Proposition 2.1.
Let $\hat{X}_{\mu}$ be the Hamiltonian vector field of the rescaled CR3BP $\hat{H}_{\mu}$. For any $\mu_{0}\in(0,1]$, the family $\mu\mapsto\hat{X}_{\mu}$ defines a one-parameter analytic perturbation (in $\nu=\mu^{1/3}$) of the Hill’s problem $\hat{X}_{0}$, in the region $q_{1}>-\mu_{0}^{-1/3}$.
@@
Figure 2. Energy levels of the five Lagrange points in the standard CR3BP $H_{\mu}$ (left) and in the rescaled CR3BP $\hat{H}_{\mu}$ (right) as a function of the mass ratio $\mu\in[0,1]$. In the limit $\mu\to 0$, the rescaled Hamiltonian corresponds to the Hill’s problem $\hat{H}_{0}$, in which only the two collinear Lagrange points $L_{1}$ and $L_{2}$ remain, both at the same energy level.
Remark 2.2.
@@
2.4. Non-degeneracy of Periodic and Symmetric Orbits
@@
Definition 2.3.
A periodic orbit $\gamma$ of a Hamiltonian system is said to be non-degenerate if the linearized return map satisfies $\det(d\psi-\operatorname{Id})\neq 0$, i.e., the differential $d\psi$ has no eigenvalue equal to $1$.
@@
2.5. Perturbation of Periodic Orbits from Hill’s Problem to CR3BP
Applying Poincaré’s perturbation theory of periodic orbits, we describe how non-degenerate periodic orbits of the Hill’s problem persist in the CR3BP for small mass ratios. Closely related approach is the method of generating solutions studied in the works of Perko, Hénon, Bruno, and Varin [perko_periodic_1983, henon_generating_1997, bruno-varin_periodic_2007], which continues planar symmetric orbits of Hill’s problem to the CR3BP within the perturbative framework. Here, we state a general perturbative result (Theorem 2.7) on fixed energy surfaces, with leading-order approximations which will be used for numerical continuation of these solutions.
Theorem 2.7.
@@ 2.5. Perturbation of Periodic Orbits from Hill’s Problem to CR3BP
In the case where $\gamma_{0}$ is symmetric, Lemma 2.6 ensures that it is non-degenerate as a symmetric periodic orbit. The same implicit function theorem argument then yields a smooth family of symmetric periodic orbits $\gamma_{\mu}$. ∎
Remark 2.8.
@@
2.6. Numerical Continuation with respect to Mass Parameter
The leading-order approximations of Theorem 2.7 provide initial guesses for numerically continuing periodic orbits when the mass ratio is small. This continuation can be carried out from Hill’s problem to the CR3BP, or vice versa, or between CR3BP systems with different small mass ratios. We make the latter precise in the following corollary.
Corollary 2.9.
Consider two small mass ratios $\mu_{1}>\mu_{2}>0$. Let $\gamma_{1}$ be a periodic orbit in the CR3BP system $H_{\mu_{1}}$ with mass ratio $\mu_{1}$, with Jacobi energy $c_{1}$ and period $T_{1}$, assumed to be a continuation from Hill’s problem as in Theorem 2.7. Then there exists a corresponding periodic orbit $\gamma_{2}$ for $H_{\mu_{2}}$:
  1. (1)
@@ 2.6. Numerical Continuation with respect to Mass Parameter
and the period remains unchanged to leading order: $T=T_{\text{prev}}$. The resulting prediction is then refined using a shooting method with Newton-type correction that ensures periodicity at the updated parameters. This process is summarized in Algorithm 1. We shall use this continuation for computation of bifurcation surfaces, i.e. bifurcation graphs of halo orbits and vertical collision orbits for the range $\mu\in[0,0.5]$ in Section 4.4.
Input: Periodic orbit $(X_{\text{prev}},T_{\text{prev}},c_{\text{prev}})$ at $\mu_{\text{prev}}$, step size $\Delta\mu$
Output: Periodic orbit $(X,T,c)$ at $\mu=\mu_{\text{prev}}+\Delta\mu$
$\mu\leftarrow\mu_{\text{prev}}+\Delta\mu$;
$h\leftarrow\mu_{\text{prev}}^{-2/3}\left(c_{\text{prev}}+(1-\mu_{\text{prev}})+\frac{(1-\mu_{\text{prev}})^{2}}{2}\right)$;
$c\leftarrow\mu^{2/3}h-(1-\mu)-\frac{(1-\mu)^{2}}{2}$;
$X\leftarrow m_{\mu}+(X_{\text{prev}}-m_{\mu_{\text{prev}}})\cdot(\mu/\mu_{\text{prev}})^{1/3}$;
$T\leftarrow T_{\text{prev}}$;
Correct $(X,T)$ using differential correction at fixed Jacobi energy $c$;
return $(X,T,c)$;
@@
until convergence, where a vector $X_{i}$ is considered to be a periodic orbit if $|F(X_{i})|<10^{-10}$. The computation of Jacobian $DF(X_{i})$ involves the linearized flow $d\varphi^{T}$, which is obtained by integrating the first-order variational equations along the trajectory.
3. Moser Regularization
@@ 3.1. Moser’s Transformation
Figure 3. Schematic picture of Moser regularization, where the momentum curve (also known as a hodograph) of an orbit is mapped to a unit sphere. The figure shows the case for the planar problem, where the unregularized momentum coordinates $(x_{1},x_{2})$ are mapped to $(\xi_{0},\xi_{1},\xi_{2})$ on the sphere via inverse stereographic projection, with $\xi_{0}$ representing the vertical axis. Collision orbits correspond to trajectories passing through the north pole $\xi_{0}=1$, and the position coordinates correspond to tangent vectors at each point.
3.2. Moser Regularization of CR3BP and Hill’s Problem
@@
at the singularity at $q=0$, where $g\neq 0$ is a constant and $\tilde{V}$ is analytic in a neighborhood of $q=0$. After an appropriate coordinate translation that places the singularity at the origin, this general form includes the CR3BP Hamiltonian $H_{\mu}$, the rescaled CR3BP Hamiltonian $\hat{H}_{\mu}$, and the Hill’s problem $\hat{H}_{0}$. The constant $g$ equals $\mu$ for the CR3BP and $1$ for Hill’s problem. For a general description of the Moser regularization in Stark-Zeeman systems, see [cieliebak_J_invariant_2017].
@@
Remark 3.3.
The vector field $X_{Q}$ refers to the intrinsic Hamiltonian vector field of $Q$ on the restricted phase space $T^{*}S^{3}$ as defined above. In contrast, $\tilde{X}_{Q}:=(\partial_{p}Q,-\partial_{q}Q)$ is the extrinsic Hamiltonian vector field of $Q$ viewed as a function on the ambient space $\mathbb{R}^{4}\times\mathbb{R}^{4}$. For numerical integration, we work in $\mathbb{R}^{8}$ and compute $X_{Q}$ from the extrinsic vector field $\tilde{X}_{Q}$ using Lemma 3.2. The regularized dynamics is then given by the ODEs
@@
3.3. Floquet Multipliers and Bifurcations
@@
If more than four eigenvalues are equal to $1$, the orbit is degenerate and bifurcations may occur. Let $X_{s}$ be a smooth family of vector fields, and let $s\mapsto\gamma_{s}$ be a smooth one-parameter family of periodic orbits of $X_{s}$. The linearized return map $R_{s}$ then depends smoothly on $s$, and its eigenvalues can be chosen to vary continuously with $s$. In Hamiltonian systems, these eigenvalues lie on the real axis $\mathbb{R}$ or on the unit circle $S^{1}\subset\mathbb{C}$. Hence, changes in stability may occur when a pair of eigenvalues merges at $1$ (tangent bifurcation), at 1-1 (period-doubling bifurcation), or when two pairs collide on $S^{1}$ and leave the unit circle as a complex quadruple (Krein collision[howard_mackay_linear_stability_1987]. Below we describe the local bifurcations relevant to our study, following the terminology of [campbell_bifurcations_1999].
Definition 3.5 (Local Bifurcations of Periodic Orbits).
@@
Note that the modified secondary Hopf bifurcation does not involve a change in stability.
3.4. Differential Correction and Continuation

### v2

@@ Abstract
keywords:
Halo orbits, Moser regularization, Restricted three-body problem, Periodic orbits, Bifurcations
1 Introduction
@@
A detailed numerical study of near-collision periodic orbits requires a regularization framework compatible with high-precision continuation. The Kustaanheimo-Stiefel (KS) regularization [Kustaanheimo_stiefel_regularization_1965], as described for the numerical setting in [howell_breakwell_almost_rectilinear_halo_1984], has traditionally been used for this purpose. In this study, we adopt an alternative approach based on Moser’s regularization. This framework offers several advantages: it is applicable in arbitrary dimensions (e.g. planar or spatial CR3BP), preserves periodicity of orbits between the regularized and unregularized systems, and provides a direct correspondence of Floquet multipliers (Lemma 8), simplifying stability analysis. We provide a first, complete implementation of Moser regularization for numerical continuation and stability analysis of periodic orbits. Our implementation treats the dynamics as a Hamiltonian system with constraints, whose framework we review in Appendix B. In particular, the regularized phase space appears as a $6$-dimensional constrained submanifold embedded in an $8$-dimensional Euclidean space.
@@
Table 1: Summary of the orbit families studied, together with their symmetries and Conley-Zehnder indices. Symmetric orbits are symmetric with respect to $y=0$, while those marked with $x$ are symmetric with respect to $x=0$.
@@
$L_{1}$ halo (collisional) small $\mu>0$ symmetric 3, 2 4.2, 4.3
@@
Saturn-Enceladus, Earth-Moon symmetric 10, 11, 12
tri-fly (“up”, “down”) Hill’s problem symmetric ($x$) 10, 11
Saturn-Enceladus, Earth-Moon non-symmetric 10, 11
tri-fly (“spin”) Earth-Moon non-symmetric -
W4/W5 Hill’s problem symmetric ($x$) 3 4.5
@@
[…]
Notably, we describe the near-collision bifurcation structure of four families of three-lobed polar orbits, which we refer to as tri-fly orbits due to their shape. In Hill’s problem, these families are associated with a period-tripling bifurcation of the vertical collision orbit, while in the Saturn-Enceladus and Earth-Moon systems they are organized by touch-and-go and pitchfork bifurcations which occur near the triple cover of the $L_{2}$ halo orbit. Determining the bifurcation mechanism involved in these orbits for small mass parameters, such as in Saturn-Enceladus, requires both high precision and regularization. Members of these families include low-altitude south pole flyby trajectories in the Enceladus system (see Figure 25).
For detailed analysis, we compute the Conley-Zehnder indices of the periodic orbit families in this study; the Conley-Zehnder index is a symplectic invariant that remains constant along non-degenerate continuous families and whose change implies bifurcations. This index provides both a diagnostic tool for numerical detection of bifurcations and a topological classification of families, and has been recently used for numerical bifurcation studies of CR3BP and related systems [moreno_aydin_van_Koert_Frauenfelder_Koh_Bifurcation_Graphs_2024, aydin_Batkhin_hill_2025, aydin_dro_2025, kumar_moreno_networks_2025]. We give details and computation methods in Appendix D.
The paper is organized as follows. The main text focuses on the underlying celestial mechanics, while the contents of a purely mathematical nature is collected in the appendix. A brief outline is given below. Section 2 introduces the rescaled Hamiltonian formulation for CR3BP and its limiting relation to Hill’s problem. Section 3 presents the numerical framework for Moser regularization and introduces the vertical collision orbit. Section 4 presents the numerical results. In Appendix B, we provide details on Hamiltonians with constraints and the restricted phase space. We follow that in Appendix C with additional details of differential correction in Moser regularization. Appendix D discusses computations of Conley-Zehnder indices. Finally, Appendix E provides numerical data for the orbit families discussed in this work.
2 Dynamical Models and Symplectic Scaling Framework
2.1 The Circular Restricted Three-Body Problem (CR3BP)
@@
Table 2: Table of mass ratio parameters for different three-body systems.
@@
[…]
2.2 Hill’s Problem and Symmetries
@@
Figure 1: Configuration of the Lagrange points and reflection symmetries in the CR3BP (left) and Hill’s problem (right). In both models, the reflection symmetry $r_{y}$ across the plane $y=0$ is present, while the Hill’s problem exhibits additional symmetry $r_{x}$ across the plane $x=0$. Only $L_{1}$ and $L_{2}$ Lagrange points remain in Hill’s problem.
2.3 Rescaled CR3BP and the Hill’s Limit
@@
Proposition 1.
Let $\hat{X}_{\mu}$ be the Hamiltonian vector field of the rescaled CR3BP $\hat{H}_{\mu}$. For any $\mu_{0}\in(0,1]$, the family $\mu\mapsto\hat{X}_{\mu}$, $\mu\in[0,\mu_{0})$ defines a one-parameter analytic perturbation (in $\nu=\mu^{1/3}$) of the Hill’s problem $\hat{X}_{0}$, in the region $q_{1}>-\mu_{0}^{-1/3}$.
@@
Figure 2: Energy levels of the five Lagrange points in the standard CR3BP $H_{\mu}$ (left) and in the rescaled CR3BP $\hat{H}_{\mu}$ (right) as a function of the mass ratio $\mu\in[0,1]$. In the limit $\mu\to 0$, the rescaled Hamiltonian corresponds to the Hill’s problem $\hat{H}_{0}$, in which only the two collinear Lagrange points $L_{1}$ and $L_{2}$ remain, both at the same energy level.
Remark 1.
@@
2.4 Non-degeneracy of Periodic and Symmetric Orbits
@@
Definition 1.
A periodic orbit $\gamma$ of a Hamiltonian system is said to be non-degenerate if the linearized return map satisfies $\det(d\psi-\mathrm{id})\neq 0$, i.e., the differential $d\psi$ has no eigenvalue equal to $1$.
@@
2.5 Perturbation of Periodic Orbits from Hill’s Problem to CR3BP
Applying Poincaré’s perturbation theory of periodic orbits, we describe how non-degenerate periodic orbits of the Hill’s problem persist in the CR3BP for small mass ratios. Closely related approach is the method of generating solutions studied in the works of Perko, Hénon, Bruno, and Varin [perko_periodic_1983, henon_generating_1997, bruno-varin_periodic_2007], which continues planar symmetric orbits of Hill’s problem to the CR3BP within the perturbative framework. Here, we state a general perturbative result (Theorem 4) on fixed energy surfaces, with leading-order approximations which will be used for numerical continuation of these solutions.
Theorem 4.
@@ 2.5 Perturbation of Periodic Orbits from Hill’s Problem to CR3BP
In the case where $\gamma_{0}$ is symmetric, Lemma 3 ensures that it is non-degenerate as a symmetric periodic orbit. The same implicit function theorem argument then yields a smooth family of symmetric periodic orbits $\gamma_{\mu}$. ∎
Remark 2.
@@
2.6 Numerical Continuation with respect to Mass Parameter
The leading-order approximations of Theorem 4 provide initial guesses for numerically continuing periodic orbits when the mass ratio is small. This continuation can be carried out from Hill’s problem to the CR3BP, or vice versa, or between CR3BP systems with different small mass ratios. We make the latter precise in the following corollary.
Corollary 5.
Consider two small mass ratios $\mu_{1}>\mu_{2}>0$. Let $\gamma_{1}$ be a periodic orbit in the CR3BP system $H_{\mu_{1}}$ with mass ratio $\mu_{1}$, with Jacobi energy $c_{1}$ and period $T_{1}$, assumed to be a continuation from Hill’s problem as in Theorem 4. Then there exists a corresponding periodic orbit $\gamma_{2}$ for $H_{\mu_{2}}$:
  1. 1.
@@ 2.6 Numerical Continuation with respect to Mass Parameter
and the period remains unchanged to leading order: $T=T_{\text{prev}}$. The resulting prediction is then refined using a shooting method with Newton-type correction that ensures periodicity at the updated parameters. This process is summarized in Algorithm 1. We use this continuation procedure at several stages of our computations to transfer known periodic orbits from the CR3BP or Hill’s problem to different mass ratios, providing initial data for orbit continuation. It is also used in the computation of bifurcation surfaces, in particular the bifurcation graphs of halo and vertical collision orbits over the range $\mu\in[0,0.5]$ in Section 4.4.
@@
1:Periodic orbit $(X_{\text{prev}},T_{\text{prev}},c_{\text{prev}})$ at $\mu_{\text{prev}}$, step size $\Delta\mu$
2:Periodic orbit $(X,T,c)$ at $\mu=\mu_{\text{prev}}+\Delta\mu$
[…]
8:Correct $(X,T)$ using differential correction at fixed Jacobi energy $c$;
9:return $(X,T,c)$
@@
until convergence, where a vector $X_{i}$ is considered to be a periodic orbit if $|F(X_{i})|$ is sufficiently small. The computation of Jacobian $DF(X_{i})$ involves the linearized flow $d\varphi^{T}$, which is obtained by integrating the first-order variational equations along the trajectory.
3 Moser Regularization
@@ 3.1 Moser’s Transformation
Figure 3: Schematic picture of Moser regularization, where the momentum curve (also known as a hodograph) of an orbit is mapped to a unit sphere. The figure shows the case for the planar problem, where the unregularized momentum coordinates $(x_{1},x_{2})$ are mapped to $(\xi_{0},\xi_{1},\xi_{2})$ on the sphere via inverse stereographic projection, with $\xi_{0}$ representing the vertical axis. Collision orbits correspond to trajectories passing through the north pole $\xi_{0}=1$, and the position coordinates correspond to tangent vectors at each point.
3.2 Moser Regularization of CR3BP and Hill’s Problem
@@
at the singularity at $q=0$, where $g\neq 0$ is a constant and $\tilde{V}$ is analytic in a neighborhood of $q=0$. For the CR3BP and Hill’s problem, the corresponding values of $g$ and $\tilde{V}$ are given by
$\text{CR3BP:}\quad g$ $=\mu,\quad\tilde{V}(q)=-\frac{1-\mu}{\sqrt{(q_{1}+1)^{2}+q_{2}^{2}+q_{3}^{2}}}-(1-\mu)q_{1}-\frac{(1-\mu)^{2}}{2},$
$\text{Hill's problem:}\quad g$ $=1,\quad\tilde{V}(q)=\frac{1}{2}|q|^{2}-\frac{3}{2}q_{1}^{2}.$
In the CR3BP case, the Hamiltonian has been translated by
$q_{1}\leftarrow q_{1}-(1-\mu),\qquad p_{2}\leftarrow p_{2}-(1-\mu),$
so that the collision singularity is located at the origin. For a general description of the Moser regularization in Stark-Zeeman systems, see [cieliebak_J_invariant_2017].
@@
Remark 3.
The vector field $X_{Q}$ refers to the intrinsic Hamiltonian vector field of $Q$ on the restricted phase space $T^{*}S^{3}$ as defined above. In contrast, $\tilde{X}_{Q}:=(\partial_{p}Q,-\partial_{q}Q)$ is the extrinsic Hamiltonian vector field of $Q$ viewed as a function on the ambient space $\mathbb{R}^{4}\times\mathbb{R}^{4}$. For numerical integration, we work in $\mathbb{R}^{8}$ and compute $X_{Q}$ from the extrinsic vector field $\tilde{X}_{Q}$ using Lemma 7. The regularized dynamics is then given by the ODEs
@@
3.3 Floquet Multipliers and Bifurcations
@@
If more than four eigenvalues are equal to $1$, the orbit is degenerate and bifurcations may occur. Let $X_{s}$ be a smooth family of vector fields, and let $s\mapsto\gamma_{s}$ be a smooth one-parameter family of periodic orbits of $X_{s}$. The linearized return map $R_{s}$ then depends smoothly on $s$, and its eigenvalues can be chosen to vary continuously with $s$: the latter claim follows from II, Theorem 5.1 and 5.2 in [kato_eigenvalues]. In Hamiltonian systems, these eigenvalues lie on the real axis $\mathbb{R}$ or on the unit circle $S^{1}\subset\mathbb{C}$. Hence, changes in stability may occur when a pair of eigenvalues merges at $1$ (tangent bifurcation), at 1-1 (period-doubling bifurcation), or when two pairs collide on $S^{1}$ and leave the unit circle as a complex quadruple (Krein collision[howard_mackay_linear_stability_1987]. Below we describe the local bifurcations relevant to our study, following the terminology of [campbell_bifurcations_1999].
Definition 3 (Local Bifurcations of Periodic Orbits).
@@
See Figure 4 for a visual description. Note that the modified secondary Hopf bifurcation does not involve a change in stability.
$\gamma_{s_{0}-\varepsilon}$
$\gamma_{s_{0}}$
$\gamma_{s_{0}+\varepsilon}$
$\gamma^{\prime}$
$\gamma_{s_{0}-\varepsilon}$
$\gamma_{s_{0}}$
$\gamma_{s_{0}+\varepsilon}$
$\gamma_{s_{0}-\varepsilon}$
$\gamma_{s_{0}}$
$\gamma_{s_{0}+\varepsilon}$
$\gamma_{s_{0}-\varepsilon}$
$\gamma_{s_{0}}$
$\gamma^{\prime}_{s_{0}-\varepsilon}$
Tangent
Period doubling
Fold
Krein
Figure 4: A schematic picture of the behavior of the eigenvalues during a bifurcation, from left to right: a tangent bifurcation, a fold bifurcation, a period doubling bifurcation, and a secondary Hopf bifurcation. The bifurcation parameter $s$ is on the vertical axis. Note that there are twice as many eigenvalues due to a symmetry property: if $\lambda$ is an eigenvalue, then so are $\bar{\lambda}$, $1/\lambda$ and $1/\bar{\lambda}$.
3.4 Differential Correction and Continuation